OpenStudy (mrdoodbromanguy):

A bag contains 3 red marbles, 5 yellow marbles, and 4 blue marbles. One marble is chosen at random. What is the probability that the chosen marble will be blue? A. 1/12 B. 1/4 C. 1/3 D. 3/4

OpenStudy (gm100299):

3+5+4 = 12 (the total number of marbles) 4 blue marbles = 4/12 reduce = 1/3

OpenStudy (mrdoodbromanguy):

thx dood :)